OpenStudy (anonymous):

Solve 4x + 3y = 13x + 5y = 33 (2, 7) (-2, 7) (2, -7) (-2, -7)

OpenStudy (anonymous):

4x + 3y = 13x + 5y = 33 these systems you write are written non-standard way.... is the system supposed to look like this: 4x + 3y = 13 x + 5y = 33 ???

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

That's the way I was TAUGHT, @dpaInc

OpenStudy (anonymous):

So rude....

OpenStudy (anonymous):

@dpaInc u r genius

OpenStudy (anonymous):

excuse me.... how was i rude? i'm just clarifying your question....

OpenStudy (anonymous):

My error

OpenStudy (anonymous):

It came off as negative, my mistake

OpenStudy (anonymous):

You were correct

OpenStudy (anonymous):

According to the choices, the equations are 4x+3y=13 and x+5y=33...the ans is (-2,7) Nothing can be done @InsanelyChaotic ....

OpenStudy (anonymous):

Your professor is the one to blame...You should surely tell him all this..

OpenStudy (anonymous):

I'm so applaud by all of this, not you, but my professor

OpenStudy (anonymous):

The lack of education

OpenStudy (anonymous):

Even when I find out an accurate answer, it doesn't even apply

OpenStudy (anonymous):

Don't get depressed..May be you will come up with correct equations tomorrow if your professor corrects himself.